Please Help. I believe that A+B ARE DONE , BUT C IS GIVING ME A PROBLEM. i HAVE WORKED THIS PROBLEM SEVERAL TIMES AND CAN'T FIGURE IT OUT??

Scenario #1: You handle internal accounting for Triad Homes, a land/housing developer. The company purchased land in South Florida and developed 250 acres. Two orders were recently placed with a nursery by the company’s landscapers. The invoices from the orders do not list the per-item price.
a) After contacting the Nursery, they inform you that the first order that was placed was for 12 Indian Hawthorn bushes and 5 Maple trees at a cost of $300.00. Use x to represent bushes and y to represent the trees. Write an equation that illustrates the situation.
12X+5Y=300 <--- IS THIS CORRECT????
b) The Nursery later indicates that the second order called for 6 Indian Hawthorn bushes and 8 Maple trees at a cost of $315.00. Write an equation that illustrates this situation using the same variables you used in part a.
6X+8Y=315 <----- IS THIS CORRECT?
c) Use the equations from parts a and b as a system of equations. Use elimination to determine the cost of each. Describe the steps you used to solve the problem

Answer by stanbon(75887)

a) After contacting the Nursery, they inform you that the first order that was placed was for 12 Indian Hawthorn bushes and 5 Maple trees at a cost of $300.00. Use x to represent bushes and y to represent the trees. Write an equation that illustrates the situation.
12X+5Y=300 <--- IS THIS CORRECT????
Yes, correct
--------------------------
b) The Nursery later indicates that the second order called for 6 Indian Hawthorn bushes and 8 Maple trees at a cost of $315.00. Write an equation that illustrates this situation using the same variables you used in part a.
6X+8Y=315 <----- IS THIS CORRECT?
Yes, correct
--------------------------
c) Use the equations from parts a and b as a system of equations. Use elimination to determine the cost of each. Describe the steps you used to solve the problem.
-----
System:
12x + 5y = 300
6x + 8y = 315
----------------
Multiply thru the 2nd equation by 2 to get the following system:
12x + 5y = 300
12x + 16y = 630
-----------------
Subtract the 1st equation from the 2nd and solve for "y":
11y = 330
y = 30 (# of Maple trees)
-----------------------------
Substitute into 6x + 8y = 315 and solve for "x":
6x + 8(30) = 315
6x + 240 = 315
6x = 75
x = 12.5 (# of bushes bought)
==============================
Cheers,
Stan H.
==============================